1) How many MOLECULES of boron trichloride are present in 4.14 moles of this compound ?

2) How many MOLES of boron trichloride are present in 1.34×1022 molecules of this compound ?

3) How many ATOMS of sulfur are present in 3.55 moles of sulfur dichloride ?

4)  How many MOLES of chlorine are present in 3.07×1022 molecules of sulfur dichloride ?
Answer #1

Ans1 Step:1 get the chemical formula of compounds Boron trichloride= B​​​​​Cl​​​​3 this compound contain one Boton atom and 3 chlorine atom

Step 2 get the atomic weight. Of each element :- atomic weight of boron :- 10 atomic weight of Chloe :- 35

Step 3 molecular weight of the compound boron trichloride :- 10+3×35 = 117g/mol

Step 4 calculate the number of mole's or given number of mole's is 4.14 moles

Step 5 multiple moles by Avogadro's number 4.14 moles ×6.022×10^23= 24.328 molecule's Ans

Ans 2

Formula used :- number of moles = number of molecule's ÷ avogardo number

n = N÷NA  

= 1.34× 10^22÷6.022×10^23=2.2251moles

And 3 chemical formula of sulfur dichloride

SCl2

1 moles of sulfur dichloride = 1mole of sulfur + 2 mole of dichloride

3..55 moles of sulfur dichloride = 1mole of sulfur

3.55 ×6.022×10^23 = 21.3781×10^23 atom

If u satisfied with my answer then hit the like below and for equiry comment please